# On the Identification of Latent Objectives in Stochastic Control

### v1

@@ I Introduction
(I.2) $dX_{t}=b(t,X_{t},u_{t})dt+\sigma(t,X_{t},u_{t})dB_{t}.$
The control processes are assumed to take values in a closed set $U$ of $\mathbb{R}^{d_{1}}$. The precise definition of the class $\mathcal{U}$ of control processes is given in the Section II below.
Suppose that the probability law $\mu_{t}$ of some state-control process $(X_{t}^{*},u_{t}^{*})$ is given for each time $t\in[0,T]$. Then the problem is to find suitable objective functions $f$ and $g$ such that $\{u_{t}^{*}\}$ is optimal for (I.1).
Let $\mathcal{D}$ be a given set of $(f,g)$ of functions such that $f$ is Borel measurable function on $0,T]\times\mathbb{R}^{d}\times U$ and $g$ is Borel measurable function on $\mathbb{R}^{d}$. For $(f,g)\in\mathcal{D}$, the functions $f$ and $g$ are interpreted as latent running cost and terminal cost functions in the forward optimal control problem (I.1), respectively. Then, to analyze the inverse problem, we introduce the functional
$V(f,g):$ $=\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(t,x,u)\mu_{t}(dxdu)dt$
$\quad+\int_{\mathbb{R}^{d}}g(x)\tilde{\mu}_{T}(dx)-J^{*}(f,g)$
where $\tilde{\mu}_{t}$ denotes the probability law of $X_{t}^{*}$ and
(I.3) $J^{*}(f,g)=\inf_{u\in\mathcal{U}}J(u;f,g).$
If $X_{t}^{*}$ is governed by the controlled SDE above, i.e., if $X_{t}^{*}=X_{t}^{u^{*}}$ then
@@
for any $(f,g)\in\mathcal{D}$. Moreover, if $u^{*}$ is optimal to the forward control problem (I.1) for some $(f^{*},g^{*})\in\mathcal{D}$, then
@@
The main contribution of the paper is a novel connection between this inverse control formulation and a generalized dynamic Schrödinger problem. For the Schrödinger problem we refer to, e.g, Schrödinger [21] and [22], Bernstein [3], Jamison [8, 9], Dai Pra [5], Mikami [12], Nagasawa [16], Nelson [18], Zambrini [24], and the works cited therein. We show that the optimal value of the inverse problem coincides with that of an stochastic control problem with relaxed marginal constraints. This result establishes a variational link between inverse stochastic control and stochastic optimal transport, providing a new conceptual perspective on inverse inference in controlled diffusions. We also establish existence of optimal solutions to the inverse problem under mild assumptions and discuss additional structural characterizations.
@@ II Existence
Under (A1), for any given $u\in\mathcal{U}$, there exists a unique solution $\{X_{t}^{u}\}_{0\leq t\leq T}$ of the controlled stochastic differential equation (I.2) with initial condition $X_{0}$ such that
(II.1) $\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{u}|^{2}\right]<\infty.$
Denote by $P_{t}^{u}$ the law of $(X_{t}^{u},u_{t})$ for each $t\in[0,T]$, i.e., $P_{t}^{u}=\mathbb{P}(X_{t}^{u},u_{t})^{-1}$. Also denote by $\tilde{P}_{t}^{u}$ the law of $X_{t}^{u}$, i.e., $\tilde{P}_{t}^{u}(A)=P_{t}^{u}(A\times U)$, $A\in\mathcal{B}(\mathbb{R}^{d})$. With these notation,
@@
Proposition II.1.
Let $(A1)$ hold. Suppose that $\mu_{t}$ is the marginal law of $(X_{t}^{u^{*}},u^{*}_{t})$ for each $t\in[0,T]$, where $u^{*}=\{u^{*}_{t}\}_{0\leq t\leq T}\in\mathcal{U}$ is an optimal policy to the problem (I.1) for some $(f^{*},g^{*})\in\mathcal{D}$. Then, this $(f^{*},g^{*})$ is a minimizer for the inverse problem (I.4) and we have
@@
Under (A1)–(A3), by (II.1) it is straightforward to check that $V^{*}=\inf_{(f,g)\in\mathcal{D}}V(f,g)>$-$\infty$.
@@
    The class $\{\phi_{f,g};(f,g)\in\mathcal{D}\}$ is contained in $C([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$ and is equi-uniformly continuous.
Example II.2.
@@
Proposition II.3.
Suppose that (A1)(A1)$(A5)$ hold. Then there exists a minimizer of the inverse problem (I.4).
@@
$\varepsilon_{k}$ $:=\sup_{u\in\mathcal{U}}\mathbb{E}\int_{0}^{T}|f_{n_{k}}(t,X_{t}^{u},u_{t})-f^{*}(t,X_{t},u_{t})|\,dt$
$\qquad+\mathbb{E}|g_{n_{k}}(X_{T}^{u})-g^{*}(X_{T}^{u})|\to 0,$
@@
Example II.4.
We consider a one-dimensional stochastic control system governed by the equation
$dX^{u}_{t}=u_{t}\,dt+\frac{1}{10}\,dB_{t},$
where $X_{0}$ is drawn from the standard normal distribution, and $B$ is a standard Brownian motion. The objective is to minimize the cost functional
$J(u;f,g)=\mathbb{E}\int_{0}^{1}\left(10|X^{u}_{t}|^{2}+\theta|u_{t}|^{2}\right)dt,$
where the admissible control space is $U=\mathbb{R}$, and $\theta>0$ is a parameter. This is a special instance of the linear quadratic control framework, for which the value function $v$ admits a closed-form expression. To evaluate the effectiveness of our method, we simulate $10^{4}$ sample trajectories $(X^{(j)}_{t_{i}},u^{(j)}_{t_{i}})$ for $j=1,\ldots,10000$ and $t_{i}=i/1000$ with $i=0,\ldots,1000$, using the above analytical formula for the optimal control with the true parameter $\theta=1$. These synthetic samples are treated as observed data. This example is taken from [17], where the true parameter is estimated by solving a root-finding problem. Here, instead, we frame the estimation of $\theta$ as a stochastic optimization task. We minimize the empirical approximation of the cost functional,
$\frac{1}{10000}\sum_{j=1}^{10000}\sum_{i=0}^{999}\left(10|X^{(j)}_{t_{i}}|^{2}+\theta|u^{(j)}_{t_{i}}|^{2}\right)\Delta t-v(0,X^{(j)}_{0};\theta),$
over $\theta\in[0.0001,10]$, using the Adam optimizer with mini-batch updates. Over 100 independent runs, the average of the estimated values of $\theta$ is found to be $0.9976$, with a standard deviation of $6.0583\times 10^{-3}$.
@@
III Connection with stochastic optimal transport
In this section, we give a connection between (I.4) and stochastic optimal transport such as the Schrödinger problem.
@@ III Connection with stochastic optimal transport
    where $f_{0}\in C_{b}([0,T]\times\mathbb{R}^{d}\times U)$, $g_{0}\in C_{b}(\mathbb{R}^{d})$, and the class $\{\phi_{f,g};(f,g)\in\tilde{\mathcal{D}}\}$ is non-empty, convex, symmetric at origin, and compact in $C_{b}([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$.
Example III.1.
A class of neural networks satisfies (B1).
@@
$\rho(\pi,\nu)$ $=\sup_{(f,g)\in\tilde{\mathcal{D}}}\bigg\{\int_{0}^{T}\langle f(t,\cdot,\cdot),\pi_{t}-\nu_{t}\rangle\,dt$
$\quad+\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle\bigg\}.$
@@
(III.1) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u;f_{0},g_{0})<\infty$
and that for any $(f,g)\in\mathcal{D}$ there exists a minimizer $u_{f,g}\in\mathcal{U}$ for the forward control problem (I.1). Then, $V^{*}$ is finite and
$V^{*}$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),\mu_{t}\rangle\,dt+\langle g_{0},\tilde{\mu}_{T}\rangle$
$\quad-\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u;f_{0},g_{0}).$
Proof.
Step (i). Take an arbitrary element $(f_{1},g_{1})\in\tilde{\mathcal{D}}$ and $\lambda>0$. Since $f_{0}$, $f_{1}$, $g_{0}$, and $g_{1}$ are all bounded, we have $J^{*}(f_{0}+\lambda f_{1},g_{0}+\lambda g_{1})>$-$\infty$ and so $V^{*}<\infty$.
Let $u_{0}\in\mathcal{U}$ be such that $\rho(P^{u},\mu)=0$. Then, since $\tilde{D}$ is symmetric at origin, for any $(f,g)\in\tilde{\mathcal{D}}$,
@@
$\int_{0}^{T}\langle f_{0}+\lambda f,\mu_{t}\rangle\,dt+\langle g_{0}+\lambda g,\tilde{\mu}_{T}\rangle$
$-J^{*}(f_{0}+\lambda f,g_{0}+\lambda g)$
@@
is concavelike in $\tilde{\mathcal{D}}$, i.e., for any $(f_{1},g_{1}),(f_{2},g_{2})\in\tilde{\mathcal{D}}$, $\alpha\in[0,1]$, there exists $(f,g)\in\tilde{\mathcal{D}}$ such that $\alpha\Psi[u,(f_{1},g_{1})]+(1-\alpha)\Psi[u,(f_{2},g_{2})]\leq\Psi[u,(f,g)]$ for any $u\in\mathcal{U}$. Moreover, the function $\Psi[u,(f,g)]$ is convexlike in $\mathcal{U}$, i.e., for any $u,\tilde{u}\in\mathcal{U}$, $\alpha\in[0,1]$, there exists $u^{\prime}\in\mathcal{U}$ such that $\alpha\Psi[u,(f,g)]+(1-\alpha)\Psi[\tilde{u},(f,g)]\geq\Psi[u^{\prime},(f,g)]$ for any $(f,g)\in\tilde{\mathcal{D}}$. Indeed, we can always take an optimal $u_{f,g}$ for $u^{\prime}$. Also, by (B1), the set $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is compact. Hence, applying Sion’s minimax theorem [23, Theorem 4.1], we get
@@
$-V^{*}$ $=\sup_{\lambda>0}\sup_{(f,g)\in\tilde{\mathcal{D}}}\inf_{u\in\mathcal{U}}\left\{J(u;f_{0},g_{0})+\lambda\Psi[u,(f,g)]\right\}$
$=\sup_{\lambda>0}\inf_{u\in\mathcal{U}}\hat{J}_{\lambda}(u)$
@@
$\hat{J}_{\lambda}(u)$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{u}-\mu_{t}\rangle\,dt+\langle g_{0},P_{T}^{u}-\tilde{\mu}_{T}\rangle$
$\quad+\lambda\rho(P^{u},\mu),\quad u\in\mathcal{U}.$
@@
Step (iii). Let $\varepsilon>0$ be fixed. Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a positive sequence such that $\lim_{n\to\infty}\lambda_{n}=0$ and
@@
holds for some $\delta>0$. Then there exists a subsequence $\{n_{k}\}$ such that $\lim_{k\to\infty}\lambda_{n_{k}}\rho(P^{u_{n_{k}}},\mu)=5\delta$. Our assumption in the theorem means the existence of $u^{*}\in\mathcal{U}$ such that $\rho(P^{u^{*}},\mu)=0$. Thus $\hat{J}_{\lambda_{n}}(u_{n})\leq\hat{J}(u^{*})+\varepsilon_{n}$ for any $n$. In particular, the sequence $\{\hat{J}_{\lambda_{n_{k}}}(u_{n_{k}})\}_{k=1}^{\infty}$ is bounded, whence we can take a further subsequence $\{n_{k_{m}}\}$ such that
@@
where $\bar{\lambda}_{m}=\lambda_{n_{k_{m}}}$ and $\bar{u}_{m}=u_{n_{k_{m}}}$. Then put $\bar{\gamma}_{m}=\gamma(P^{\bar{u}_{m}},\mu)$. Now choose $m_{0}$ and $m_{1}$ such that $\kappa<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta$, $\hat{J}_{\bar{\lambda}_{m_{1}}}(\bar{u}_{m_{1}})<\kappa+\delta$, $\bar{\lambda}_{m_{1}}>7\bar{\lambda}_{m_{0}}$, and that $3\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}<7\delta$, where $\bar{\varepsilon}_{m}=\varepsilon_{n_{k_{m}}}$. With these choices it follows that
$\kappa$ $<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta\leq\inf_{u\in\mathcal{U}}\hat{J}_{\bar{\lambda}_{m_{0}}}(u)+\bar{\varepsilon}_{m_{0}}+\delta$
$\leq\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{1}})+\bar{\varepsilon}_{m_{0}}+\delta$
$=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{\bar{u}_{m_{1}}}-\mu_{t}\rangle\,dt+\langle g_{0},\tilde{P}_{T}^{\bar{u}_{m_{1}}}-\tilde{\mu}_{T}\rangle$
$\quad+\frac{\bar{\lambda}_{m_{0}}}{\bar{\lambda}_{m_{1}}}\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta.$
@@
$\bar{\lambda}_{m_{0}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta<(1/7)\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta$
$<2\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}-\delta,$
@@
$\inf_{u\in\mathcal{U}}\hat{J}_{\lambda_{n}}(u)\geq\hat{J}_{\lambda}(u_{n})-\varepsilon_{n}$
$\geq\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})]+\lambda_{n}\rho(P^{u_{n}},\mu)-\varepsilon_{n}.$
@@
Here, we shall discuss a connection between Theorem III.2 and the Schrödinger bridge problems. Consider the case where the forward control problem (I.1) is described as
$J(u)=\mathbb{E}\int_{0}^{1}|u_{t}|^{2}dt,$
@@
$dX_{t}=u_{t}dt+dB_{t},\quad X_{0}\sim\tilde{\mu}_{0}.$
The class $\mathcal{D}$ of latent objectives is assumed to be given by

### v2

# Inverse Stochastic Control via Generalized Schrödinger Problems

@@ 1 Introduction
(1.2) $dX_{t}^{u}=b(t,X_{t}^{u},u_{t})dt+\sigma(t,X_{t}^{u},u_{t})dB_{t}.$
The control processes are assumed to take values in a closed set $U$ of $\mathbb{R}^{d_{1}}$. The precise definition of the class $\mathcal{U}$ of control processes is given in the Section 2 below.
Suppose that a family of probability measures $\{\mu_{t}\}_{0\leq t\leq T}$ representing the observed state-control distribution is given. Importantly, $\mu_{t}$ need not be realizable as the marginal law of any controlled process; the formulation below is meaningful for any family of measures satisfying mild integrability conditions (see Section 2). In the canonical case where $\mu_{t}$ is indeed the law of $(X_{t}^{u^{*}},u_{t}^{*})$ for some $u^{*}\in\mathcal{U}$, the problem reduces to finding objective functions $f$ and $g$ such that $u^{*}$ is optimal for (1.1).
Let $\mathcal{D}$ be a given set of $(f,g)$ of functions such that $f$ is Borel measurable function on $0,T]\times\mathbb{R}^{d}\times U$ and $g$ is Borel measurable function on $\mathbb{R}^{d}$. For $(f,g)\in\mathcal{D}$, the functions $f$ and $g$ are interpreted as latent running cost and terminal cost functions in the forward optimal control problem (1.1), respectively. To motivate the formulation, we argue informally as follows. We introduce the functional
$V(f,g):$ $=\int_{0}^{T}\int_{\mathbb{R}^{d}\times U}f(t,x,u)\mu_{t}(dxdu)dt+\int_{\mathbb{R}^{d}}g(x)\tilde{\mu}_{T}(dx)-J^{*}(f,g)$
where $\tilde{\mu}_{t}$ denotes the probability law of $X_{t}^{u^{*}}$ and
(1.3) $J^{*}(f,g)=\inf_{u\in\mathcal{U}}J(u;f,g).$
If $(X_{t}^{u^{*}},u_{t}^{*})$ satisfies the controlled SDE above for some $u^{*}\in\mathcal{U}$, then
@@
for any $(f,g)\in\mathcal{D}$. Moreover, if $u^{*}$ is optimal to the forward control problem (1.1) for some $(f^{*},g^{*})\in\mathcal{D}$, then
@@
The main contribution of the paper is a novel connection between this inverse control formulation and a generalized dynamic Schrödinger problem. For the Schrödinger problem we refer to, e.g, Schrödinger [21] and [22], Bernstein [3], Jamison [8, 9], Dai Pra [5], Mikami [12], Nagasawa [16], Nelson [18], Zambrini [24], and the works cited therein. We show that the optimal value of the inverse problem coincides with that of a stochastic control problem with relaxed marginal constraints. This result establishes a variational link between inverse stochastic control and stochastic optimal transport, providing a new conceptual perspective on inverse inference in controlled diffusions. We also establish existence of optimal solutions to the inverse problem under mild assumptions and discuss additional structural characterizations.
@@ 2 Existence
Under (A1), for any given $u\in\mathcal{U}$, there exists a unique solution $\{X_{t}^{u}\}_{0\leq t\leq T}$ of the controlled stochastic differential equation (1.2) with initial condition $X_{0}$ such that
(2.1) $\mathbb{E}\left[\sup_{0\leq t\leq T}|X_{t}^{u}|^{2}\right]<\infty.$
Denote by $P_{t}^{u}\in\mathcal{P}(\mathbb{R}^{d}\times U)$ the probability distribution of $(X_{t}^{u},u_{t})$ under $\mathbb{P}$ for each $t\in[0,T]$. Also denote by $\tilde{P}_{t}^{u}$ the law of $X_{t}^{u}$, i.e., $\tilde{P}_{t}^{u}(A)=P_{t}^{u}(A\times U)$, $A\in\mathcal{B}(\mathbb{R}^{d})$. With these notation,
@@
Proposition 2.1.
Let $(A1)$ hold. Suppose that $\mu_{t}$ is the marginal law of $(X_{t}^{u^{*}},u^{*}_{t})$ for each $t\in[0,T]$, where $u^{*}=\{u^{*}_{t}\}_{0\leq t\leq T}\in\mathcal{U}$ is an optimal policy to the problem (1.1) for some $(f^{*},g^{*})\in\mathcal{D}$. Then, this $(f^{*},g^{*})$ is a minimizer for the inverse problem (1.4) and we have
@@
Under (A1)–(A3), by (2.1) it is straightforward to check that $V^{*}=\inf_{(f,g)\in\mathcal{D}}V(f,g)>$-$\infty$.
@@
    The class $\{\phi_{f,g}:(f,g)\in\mathcal{D}\}$ is contained in $C([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$ and is equi-uniformly continuous.
Example 2.2.
@@
Proposition 2.3.
Suppose that (A1)(A1)$(A5)$ hold. Then there exists a minimizer of the inverse problem (1.4).
@@
$\varepsilon_{k}$ $:=\sup_{u\in\mathcal{U}}\mathbb{E}\int_{0}^{T}|f_{n_{k}}(t,X_{t}^{u},u_{t})-f^{*}(t,X_{t}^{u},u_{t})|\,dt+\mathbb{E}|g_{n_{k}}(X_{T}^{u})-g^{*}(X_{T}^{u})|\to 0,$
@@
Example 2.4.
Consider the one-dimensional system
$dX^{u}_{t}=u_{t}\,dt+\tfrac{1}{10}\,dB_{t},\quad X_{0}\sim N(0,1),$
with the parametric cost class $\mathcal{D}=\{(f_{\theta},0):\theta>0\}$, where $f_{\theta}(t,x,u)=10|x|^{2}+\theta|u|^{2}$. For each $\theta>0$, the value function takes the form $v(t,x;\theta)=r_{t}(\theta)x^{2}+s_{t}(\theta)$, where $r_{t}(\theta)$ satisfies the Riccati equation $\dot{r}_{t}=r_{t}^{2}/\theta-10$, $r_{1}=0$, with explicit solution
$r_{t}(\theta)=\sqrt{10\theta}\,\tanh\!\bigl(\sqrt{10/\theta}\,(1-t)\bigr).$
Since $X_{0}\sim N(0,1)$, the optimal value is
$J^{*}(f_{\theta},0)=\sqrt{10\theta}\,\tanh\!\bigl(\sqrt{10/\theta}\bigr)-\tfrac{\theta}{100}\ln\cosh\!\bigl(\sqrt{10/\theta}\bigr).$
Suppose the observed data are generated by the optimal control $u_{t}^{*}=-(r_{t}(\theta^{*})/\theta^{*})X_{t}^{u^{*}}$ under the true parameter $\theta^{*}=1$. By optimality, $V(f_{\theta^{*}},0)=0$, and the strict convexity of the LQ problem implies $V(f_{\theta},0)>0$ for all $\theta\neq\theta^{*}$. Hence $V^{*}=0$ and $\theta^{*}=1$ is the unique minimizer of $\theta\mapsto V(f_{\theta},0)$ over $\theta>0$, confirming that the inverse problem (1.4) correctly recovers the true parameter.
@@
3 Connection with stochastic optimal transport
In this section, we give a connection between (1.4) and stochastic optimal transport such as the Schrödinger problem.
@@ 3 Connection with stochastic optimal transport
    where $f_{0}\in C([0,T]\times\mathbb{R}^{d}\times U)$ with $\inf_{(t,x,u)}f_{0}(t,x,u)>$-$\infty$, $g_{0}\in C_{b}(\mathbb{R}^{d})$, and the class $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is non-empty, convex, symmetric at origin, and compact in $C_{b}([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$.
Example 3.1.
A class of neural networks satisfies (B1). More precisely, let $f_{\theta}(t,x,u)$ and $g_{\theta}(x)$ be neural networks with a fixed architecture, parameterized by weights $\theta\in\Theta$, where $\Theta\subset\mathbb{R}^{m}$ is a compact, convex, and symmetric set. If the map $\theta\mapsto\phi_{f_{\theta},g_{\theta}}$ is affine (as in the case of a single hidden layer with linear output), then the class $\{\phi_{f_{\theta},g_{\theta}}:\theta\in\Theta\}$ is a compact, convex, and symmetric subset of $C_{b}([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$, so (B1) is satisfied with $\tilde{\mathcal{D}}=\{(f_{\theta},g_{\theta}):\theta\in\Theta\}$.
@@
$\rho(\pi,\nu)$ $=\sup_{(f,g)\in\tilde{\mathcal{D}}}\bigg\{\int_{0}^{T}\langle f(t,\cdot,\cdot),\pi_{t}-\nu_{t}\rangle\,dt+\langle g,\tilde{\pi}_{T}-\tilde{\nu}_{T}\rangle\bigg\}.$
@@
(3.1) $\inf_{u\in\mathcal{U}_{\rho,0}}J(u;f_{0},g_{0})<\infty$
and that for any $(f,g)\in\mathcal{D}$ there exists a minimizer $u_{f,g}\in\mathcal{U}$ for the forward control problem (1.1). Then, $V^{*}$ is finite and
$V^{*}$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),\mu_{t}\rangle\,dt+\langle g_{0},\tilde{\mu}_{T}\rangle-\lim_{\varepsilon\searrow 0}\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}J(u;f_{0},g_{0}).$
Remark 3.3.
The assumption that a minimizer $u_{f,g}\in\mathcal{U}$ exists for every $(f,g)\in\mathcal{D}$ is used to verify the convexlike condition on $\Psi[u,(f,g)]$ required by Sion’s minimax theorem (Step (ii) of the proof below). Relaxing this to the existence of $\varepsilon$-optimal controls, or establishing it via measurable selection arguments under appropriate continuity conditions on $(f,g)\mapsto J^{*}(f,g)$, would broaden the scope of Theorem 3.2 and is a subject of future work.
Proof of Theorem 3.2.
Step (i). Take an arbitrary element $(f_{1},g_{1})\in\tilde{\mathcal{D}}$ and $\lambda>0$. Since $f_{0}$ is bounded below and $f_{1}$, $g_{0}$, $g_{1}$ are all bounded, we have $J^{*}(f_{0}+\lambda f_{1},g_{0}+\lambda g_{1})>$-$\infty$ and so $V^{*}<\infty$.
Let $u_{0}\in\mathcal{U}$ be such that $\rho(P^{u_{0}},\mu)=0$. Then, since $\tilde{\mathcal{D}}$ is symmetric at origin, for any $(f,g)\in\tilde{\mathcal{D}}$,
@@
$\int_{0}^{T}\langle f_{0}+\lambda f,\mu_{t}\rangle\,dt+\langle g_{0}+\lambda g,\tilde{\mu}_{T}\rangle-J^{*}(f_{0}+\lambda f,g_{0}+\lambda g)$
@@
is concavelike in $\tilde{\mathcal{D}}$, i.e., for any $(f_{1},g_{1}),(f_{2},g_{2})\in\tilde{\mathcal{D}}$, $\alpha\in[0,1]$, there exists $(f,g)\in\tilde{\mathcal{D}}$ such that $\alpha\Psi[u,(f_{1},g_{1})]+(1-\alpha)\Psi[u,(f_{2},g_{2})]\leq\Psi[u,(f,g)]$ for any $u\in\mathcal{U}$. Moreover, the function $\Psi[u,(f,g)]$ is convexlike in $\mathcal{U}$, i.e., for any $u,\tilde{u}\in\mathcal{U}$, $\alpha\in[0,1]$, there exists $u^{\prime}\in\mathcal{U}$ such that $\alpha\Psi[u,(f,g)]+(1-\alpha)\Psi[\tilde{u},(f,g)]\geq\Psi[u^{\prime},(f,g)]$ for any $(f,g)\in\tilde{\mathcal{D}}$. Indeed, we can always take an optimal $u_{f,g}$ for $u^{\prime}$. Note that the existence of a minimizer $u_{f,g}$ for each $(f,g)\in\mathcal{D}$, assumed in Theorem 3.2, is precisely what guarantees this convexlike condition required by Sion’s minimax theorem. Also, by (B1), the set $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is compact. Furthermore, the map $(f,g)\mapsto\Psi[u,(f,g)]$ is continuous with respect to the topology on $\tilde{\mathcal{D}}$: indeed, $\Psi[u,(f,g)]$ is linear in $(f,g)$, and since the set $\{\phi_{f,g}:(f,g)\in\tilde{\mathcal{D}}\}$ is compact in $C_{b}([0,T]\times\mathbb{R}^{d}\times U,\mathbb{R}^{2})$, it is uniformly bounded, say $\|\phi_{f,g}\|_{\infty}\leq M$ for all $(f,g)\in\tilde{\mathcal{D}}$. Hence, by the bounded convergence theorem (applicable under assumptions (A2) and (2.1)), $\Psi[u,(f,g)]$ is norm-continuous in $(f,g)$ on $\tilde{\mathcal{D}}$, and in particular upper and lower semicontinuous as required by Sion’s theorem. Hence, applying Sion’s minimax theorem [23, Theorem 4.2], we get
@@
$-V^{*}$ $=\sup_{\lambda>0}\sup_{(f,g)\in\tilde{\mathcal{D}}}\inf_{u\in\mathcal{U}}\left\{\Psi[u,(f_{0},g_{0})]+\lambda\Psi[u,(f,g)]\right\}=\sup_{\lambda>0}\inf_{u\in\mathcal{U}}\hat{J}_{\lambda}(u)$
@@
$\hat{J}_{\lambda}(u)$ $=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{u}-\mu_{t}\rangle\,dt+\langle g_{0},\tilde{P}_{T}^{u}-\tilde{\mu}_{T}\rangle+\lambda\rho(P^{u},\mu),\quad u\in\mathcal{U}.$
@@
Step (iii). Let $\varepsilon>0$ be fixed. Let $\{\lambda_{n}\}_{n=1}^{\infty}$ be a positive sequence such that $\lim_{n\to\infty}\lambda_{n}=\infty$ and
@@
holds for some $\delta>0$. Then there exists a subsequence $\{n_{k}\}$ such that $\lim_{k\to\infty}\lambda_{n_{k}}\rho(P^{u_{n_{k}}},\mu)=5\delta$. Our assumption in the theorem means the existence of $u^{*}\in\mathcal{U}$ such that $\rho(P^{u^{*}},\mu)=0$. Thus $\hat{J}_{\lambda_{n}}(u_{n})\leq\hat{J}_{\lambda_{n}}(u^{*})+\varepsilon_{n}$ for any $n$. In particular, the sequence $\{\hat{J}_{\lambda_{n_{k}}}(u_{n_{k}})\}_{k=1}^{\infty}$ is bounded, whence we can take a further subsequence $\{n_{k_{m}}\}$ such that
@@
where $\bar{\lambda}_{m}=\lambda_{n_{k_{m}}}$ and $\bar{u}_{m}=u_{n_{k_{m}}}$. Then put $\bar{\gamma}_{m}=\rho(P^{\bar{u}_{m}},\mu)$. Now choose $m_{0}$ and $m_{1}$ such that $\kappa<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta$, $\hat{J}_{\bar{\lambda}_{m_{1}}}(\bar{u}_{m_{1}})<\kappa+\delta$, $\bar{\lambda}_{m_{1}}>7\bar{\lambda}_{m_{0}}$, and that $3\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}<7\delta$, where $\bar{\varepsilon}_{m}=\varepsilon_{n_{k_{m}}}$. With these choices it follows that
$\kappa$ $<\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{0}})+\delta\leq\inf_{u\in\mathcal{U}}\hat{J}_{\bar{\lambda}_{m_{0}}}(u)+\bar{\varepsilon}_{m_{0}}+\delta\leq\hat{J}_{\bar{\lambda}_{m_{0}}}(\bar{u}_{m_{1}})+\bar{\varepsilon}_{m_{0}}+\delta$
$=\int_{0}^{T}\langle f_{0}(t,\cdot,\cdot),P_{t}^{\bar{u}_{m_{1}}}-\mu_{t}\rangle\,dt+\langle g_{0},\tilde{P}_{T}^{\bar{u}_{m_{1}}}-\tilde{\mu}_{T}\rangle+\bar{\lambda}_{m_{0}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta.$
@@
$\bar{\lambda}_{m_{0}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta<\frac{1}{7}\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}+\bar{\varepsilon}_{m_{0}}+\delta<2\delta+\bar{\varepsilon}_{m_{0}}<\bar{\lambda}_{m_{1}}\bar{\gamma}_{m_{1}}-\delta,$
@@
$\inf_{u\in\mathcal{U}}\hat{J}_{\lambda_{n}}(u)\geq\hat{J}_{\lambda_{n}}(u_{n})-\varepsilon_{n}\geq\inf_{u\in\mathcal{U}_{\rho,\varepsilon}}\Psi[u,(f_{0},g_{0})]+\lambda_{n}\rho(P^{u_{n}},\mu)-\varepsilon_{n}.$
@@
Here, we shall discuss a connection between Theorem 3.2 and the Schrödinger bridge problems. Consider the case where the forward control problem (1.1) is described as
$J(u)=\mathbb{E}\int_{0}^{T}|u_{t}|^{2}dt,$
@@
$dX_{t}^{u}=u_{t}dt+dB_{t},\quad X_{0}\sim\tilde{\mu}_{0}.$
The class $\mathcal{D}$ of latent objectives is given by